Suppose that the demand function for pizzas is Q^d = 65,800 - 1,200P and the supply function is Q^s = 4,000P - 20,000. Suppose the pizza parlor industry is effective at lobbying the government, which institutes a price floor for $15 on pizzas. Assuming that the least - cost pizza produces are the ones to produce the demanded pizzas, what is the effect on the aggregate consumer, and producer surpluses? What if instead the government raised the price to $15 using a price support program?
Answer #1

Q^d = 65,800 - 1,200P
Q^s = 4,000P - 20,000

65,800 - 1,200P = 4,000P - 20,000
5200P = 85800
P = 16.5

Price floor = 15

As the price floor is below the competitive price equilibrium, there will be no impact on the aggregate consumer, and producer surpluses. There will be no impact of the support program either.